# An Axiomatic Risk-Reward Framework for Sustainable Investing

### v3

@@ Abstract
Continued interest in sustainable investing calls for an axiomatic approach to measures of risk and reward that focus not only on financial returns, but also on measures of environmental and social sustainability, i.e. environmental, social, and governance (ESG) scores. We propose definitions for ESG-coherent risk measures and ESG reward–risk ratios based on functions of bivariate random variables that are applied to financial returns and real-time ESG scores, extending the traditional univariate measures to the ESG case. We provide examples and present an empirical analysis in which the ESG-coherent risk measures and ESG reward–risk ratios are used to rank stocks.
Keywords: Finance, ESG, risk measures, multivariate risk, sustainable investing.
@@ 1 Introduction
ESG investing encompasses a broad array of approaches, and its market practices are very heterogeneous, with different terminologies, definitions, and strategies. These practices vary due to the cultural and ideological diversity of investors (Sandberg et al.,, 2009; Widyawati,, 2020). According to Amel-Zadeh and Serafeim, (2018), the most significant motivation for incorporating ESG factors is related to financial performance, as sustainability factors are perceived as relevant to investment returns. That is, investors believe that ESG data can be used to identify potential risks and opportunities, and that such information is not yet fully incorporated into market prices. Hence, ESG information should help investors to control risk better and improve their financial performance. In line with Schanzenbach and Sitkoff, (2020) we employ the expression risk-return ESG to refer to investment strategies that use ESG factors to improve returns while lessening risk. Academic evidence on the role of ESG in enhancing performance is inconclusive. The meta-analysis conducted by Revelli and Viviani, (2015) shows that sustainable and responsible investing URI) is neither a weakness nor a strength compared with conventional investing. Similar conclusions have been drawn by Hornuf and Yüksel, (2024). [^1] We note that, if ESG information is already incorporated in market prices, any strategy that restricts investment based on ESG criteria would result in sub-optimal allocation in terms of monetary performance. Moreover we underline that risk-return ESG strategies are not necessarily more sustainable than traditional ones. Indeed, an investor may implement a contrarian-ESG approach, investing in less sustainable assets if they are expected to outperform the market.
@@
A second motivation that guides ESG strategies is the desire to improve the sustainable profile of the portfolios for ethical reasons or to improve investors’ green image (Amel-Zadeh and Serafeim,, 2018). [^2] Almel-Zadeh and Serafim reported additional motivations that push asset managers to propose ESG products, including growing client demand, the effectiveness of ESG investing in bringing about change in companies, and formal client mandates. For our purposes, these drivers can be attributed to either material or ethical motivations. Sustainability then becomes part of the investment goals, alongside monetary performance and the riskiness of the position. Schanzenbach and Sitkoff, (2020) refer to investment strategies that incorporate ESG screenings for moral or ethical reasons as collateral benefit ESG, as they aim to provide benefits to a third party, rather than to improve risk-adjusted returns. We refer to investors who include sustainability considerations for ethical reasons as ESG-oriented investors, to distinguish them from risk-return investors who care exclusively about the financial risks and returns of a position.
@@
From a theoretical point of view, risk-return ESG does not pose any specific issue, as ESG is treated as any other information (e.g. balance sheet data, macroeconomic indicators, sentiment analysis, etc.) and is integrated in the investment process without affecting the main goal of the investor: improving the risk-adjusted performance. In contrast, the ethical motivations of ESG-oriented investors challenge many of the traditional assumptions adopted in finance theory, as the inclusion of other factors in the investment process (e.g. reduction of CO2 emissions, respect of human right, exclusion of controversial sectors, etc.) are not necessarily motivated by the improvement of financial perfomance.
@@
To illustrate how the ethical motivation breaks common investing principles, consider two stocks with the same reward–risk profile, but with different ESG scores. They may be equivalent for a risk-return investor but not for an ESG-oriented investor, as one of the two companies may be more environmentally sustainable or may adopt stricter human right policies. An ESG-oriented investor may decide to deliberately worsen their risk-adjusted performance to comply with non-negotiable principles, for instance by excluding certain sectors or companies, thus reducing the diversification of their portfolio. [^3] It is common practice for sustainable investment funds to exclude tobacco, weapons, fossil fuels, and other controversial sectors. The divestment campaign aimed at South Africa’s apartheid regime in the 1980s was one of the key turning points in the history of responsible investing (Schanzenbach and Sitkoff,, 2020). The financial literature has discussed the role of sustainable investors on market equilibrium. In a seminal paper Heinkel et al., (2001) investigated the effects of exclusionary ethical investments on corporate behaviour, finding that if the percentage of ethical investors is sufficiently high, the cost of capital for polluting companies may increase. More recently, Pástor et al., (2021) built a two-factor model to study the market equilibrium in the presence of ESG investors, finding that, in equilibrium, sustainable assets have lower expected returns due to the fact that ESG investors have a preference towards these assets, and because they can be used to hedge climate risk.
In this work we develop a theoretical framework for ESG-oriented investors, aimed at measuring risk and reward incorporating non-financial considerations, supporting their decision making process. In particular, we propose axiomatic classes of measures that extend the concepts of coherent risk measures (Artzner et al.,, 1999), reward measures (Rachev et al.,, 2008), and reward-risk ratios (Cheridito and Kromer,, 2013) for ESG-oriented investors.
The measurement of risk and reward for ESG investors is strictly related to the problem of setting up optimal portfolio strategies that combine financial and sustainability goals. The problem has been addressed in the literature by extending the classical reward–risk analysis by adding a third dimension – sustainability – measured using ESG scores. ESG scores are then included in the optimal allocation problem as a constraint, or in the objective function. Such an approach has been pursued by, among others, Utz et al., (2015), Gasser et al., (2017) and Cesarone et al., (2022) who studied the efficient frontier of non-dominated portfolios in the reward-risk-ESG space, extending the traditional optimal portfolio literature that started with Markowitz, (1952), which aims to find portfolios with the highest expected return for a given level of risk, where risk is measured using the portfolio variance (Markowitz,, 1952), a coherent risk measure such as the average value at risk (AVaR)444 The AVaR is also known in the literature as Conditional Value at Risk (CVaR) or Expected Shortfall (ES) (Rockafellar et al.,, 2000), or an asymmetric deviation measure (Giacometti et al.,, 2021).555 An alternative approach for including ESG in the frontier analysis was proposed by Pedersen et al., (2021), who identified a two-dimensional ESG efficient frontier by considering the Sharpe ratio and the ESG score. One drawback of this modeling framework is that it fails to take into account the stochasticity of ESG scores: it either uses only the expected values of the scores or assumes that ESG scores can be treated as constants for the duration of the investment period.
@@
Here, we aim to address this limitation and, more generally, introduce a way to measure the risk associated with an ESG-oriented position. We propose an axiomatic approach based on the idea that risk comes from two drivers: monetary performance (i.e., the financial returns of a position) and sustainability (represented by the ESG score of the company). [^6] For convenience, in the rest of this paper, we refer to these two dimensions as the monetary and ESG components of risk. These quantities are random and not necessarily independent, and we can represent them as a bivariate random variable. Hence, it is natural to refer to the rich literature on multivariate risk measures (Jouini et al.,, 2004; Hamel,, 2009; Wei and Hu,, 2014; Ekeland et al.,, 2012). Such measures have been developed to study portfolios of non-perfectly fungible assets (e.g., assets valued in multiple currencies) or assets that are difficult to price. We propose using a bivariate risk measure to deal with a single asset that can be evaluated over two dimensions: the monetary returns and the sustainability (represented by ESG scores). We then define ESG-coherent risk measures as an extension of the coherent risk measures introduced by Artzner et al., (1999).
@@
From an empirical perspective, the analysis is motivated both by market participants’ interest in sustainability and the recent implementation of real-time ESG measures (as opposed to most ESG data that are published with annual frequency), such as the Truvalue scores issued by Factset.
Section 2 discusses our interpretation of ESG scores and outlines how we propose to use them. Section 3 introduces ESG-coherent risk measures and provides examples, while Section 4 introduces ESG reward–risk ratios. Section 5 presents an empirical application using the equity returns and real-time ESG data of the constituents of the Dow-Jones Industrial Average (DJIA) financial index. Section 6 highlights some conclusions and perspectives for future research.
@@ 2 Measuring ESG and financial performance
The first step in defining ESG risk measures is to clarify how we characterize and measure sustainability. Following an approach common in the literature, we use the ESG scores of a company as a proxy for sustainability (see, e.g., Pedersen et al.,, 2021; Utz et al.,, 2015; Cesarone et al.,, 2022). Here, we clarify how to treat such a variable from both practical and theoretical points of view, establishing the basis for measuring ESG scores in a way that allows us to consistently use them alongside monetary values for risk measurement. We emphasize that, at this stage, we keep the discussion on an abstract level, without reference to specific data providers or methodologies. This allows us not to be limited by the current state-of-the-art market ESG scores, which are still far from being standardized and comparable across data providers (see, e.g., Berg et al.,, 2022; Billio et al.,, 2021).
@@
Several approaches appear in the literature for modeling the random variable $r_{T}$ used to compute risk. Artzner et al., (1999) define coherent risk measures for a random variable that represents net worth by following the principle that “bygones are bygones,” meaning that future net worth is the only thing that matters. In an alternative approach, the random variable used to compute monetary risk represents the return of a financial position. This latter approach, which is often used in practical applications for the measurement of the risk of equity portfolios, introduces some differences in the interpretation of the axioms (see Rachev et al.,, 2011, Chapter 6). [^7] Alternatively, some authors prefer to define risk measures computed on a variable that represents losses, thus assuming that lower values of the variable are preferred by an investor. Such an assumption does not significantly alter the analysis: it simply changes some signs in the definition of risk (see, e.g., Rockafellar and Uryasev,, 2013).
@@
Similarly, we need to establish the quantity measured by the random variable ${ESG}_{T}$ and the sign convention used. In particular, one must consider whether it is a stock variable (measured at a specific point in time) or a flow variable (measured over an interval of time, as some sort of sustainability return). Based on how they are computed, we argue that ESG scores belong to the second category: indeed, they represent the current level of sustainability of the production and commercial practices of a company, which directly affects the impact of that company on the world.888 The scores are typically computed as functions of several indicators related to the production methods, the supply chain management, the industry in which the company operates, the transparency of its governance, the presence of specific policies on human rights violations, etc. We can think of ESG scores as a broad measure of externalities (positive or negative) that are generated over time. In this sense, the ESG score of a company is related to the rate at which it accumulates non-monetary “satisfaction” for the investor. The total non-monetary satisfaction for an investor is proportional to the holding time of the investment.
@@
We can then interpret the ESG scores assigned by market providers with periodicity $T$ (e.g. a day or a year), as the realizations of the random variables ${ESG}_{T}$. Assuming that the ESG scores are bounded, in the rest of the work we rescale the values of daily ESG scores such that ${ESG}_{T}\in[-1/c,1/c]$, where $T$ is equal to 1 trading day, and $c=252$. As discussed later, this standardization leads to the “typical magnitude” of ${ESG}_{T}$ for a daily time period to be comparable to the “typical magnitude” of the financial cumulative returns measured over the same time period.
We finally formalize the bivariate return of an asset as the random variable $X_{T}$ with two components: ${ESG}_{T}$ for the sustainability, and $r_{T}$ for the monetary part, computed as the cumulative log-return for a period of length $T$:
$X_{T}=\begin{bmatrix}r_{T}\\ {ESG}_{T}\end{bmatrix}$. (2)
The stochasticity of the process $\text{esg}_{t}$ can be traced to two sources: the first is the uncertainty in the measurement of its value, and the second is the uncertainty in the evolution of the sustainability policies within the company, which is driven by the choices made by the management, by market conditions, and by the regulatory framework. We also expect a correlation structure between ${ESG}_{T}$ and financial returns $r_{T}$ due to the presence of common driving factors related, for instance, to sector-wide or country-wide dynamics. With regard to a sign convention, it is natural to express ESG scores in such a way that higher scores are preferable. This is consistent with the preference for higher returns $r_{T}$.

### v4

@@ Abstract
Continued interest in sustainable investing calls for an axiomatic approach to measures of risk and reward that focus not only on financial returns, but also on measures of environmental and social sustainability, i.e. environmental, social, and governance (ESG) scores. We propose axiomatic definitions for ESG-coherent risk measures and ESG reward–risk ratios based on functions of bivariate random variables that are applied to financial returns and real-time ESG scores, extending the traditional univariate measures to the ESG case. We provide examples, discuss the dual representation, and present an empirical analysis in which the ESG-coherent risk measures and ESG reward–risk ratios are used to rank stocks.
Keywords: risk measures, multivariate risk, ESG, sustainable investing
@@ 1 Introduction
ESG investing encompasses a broad array of approaches, and its market practices are very heterogeneous, with different terminologies, definitions, and strategies. These practices vary due to the cultural and ideological diversity of investors (Sandberg et al., 2009; Widyawati, 2020). According to Amel-Zadeh and Serafeim (2018), the most significant motivation for incorporating ESG factors is related to financial performance, as sustainability factors are perceived as relevant to investment returns. That is, investors believe that ESG data can be used to identify potential risks and opportunities, and that such information is not yet fully incorporated into market prices. Hence, ESG information should help investors to control risk better and improve their financial performance. In line with Schanzenbach and Sitkoff (2020) we employ the expression risk-return ESG to refer to investment strategies that use ESG factors to improve returns while lessening risk. Academic evidence on the role of ESG in enhancing performance is inconclusive. The meta-analysis conducted by Revelli and Viviani (2015) shows that sustainable and responsible investing (SRI) is neither a weakness nor a strength compared with conventional investing. In contrast, Friede et al. (2015) combining the findings of approximately 2200 studies from 1970 to 2014 find that large majority of of the considered literature reports a positive relationship between ESG and corporate financial performance. More recently, Hornuf and Yüksel (2024) found that responsible investment neither outperforms nor underperforms the market portfolio. [^1] We note that, if ESG information is already incorporated in market prices, any strategy that restricts investment based on ESG criteria would result in sub-optimal allocation in terms of monetary performance. Moreover we underline that risk-return ESG strategies are not necessarily more sustainable than traditional ones. Indeed, an investor may implement a contrarian-ESG approach, investing in less sustainable assets if they are expected to outperform the market.
@@
A second motivation that guides ESG strategies is the desire to improve the sustainable profile of the portfolios for ethical reasons or to improve investors’ green image (Amel-Zadeh and Serafeim, 2018). [^2] Almel-Zadeh and Serafim reported additional motivations that push asset managers to propose ESG products, including growing client demand, the effectiveness of ESG investing in bringing about change in companies, and formal client mandates. For our purposes, these drivers can be attributed to either material or ethical motivations. Sustainability then becomes part of the investment goals, alongside monetary performance and the riskiness of the position. Schanzenbach and Sitkoff (2020) refer to investment strategies that incorporate ESG screenings for moral or ethical reasons as collateral benefit ESG, as they aim to provide benefits to a third party, rather than to improve risk-adjusted returns. We refer to investors who include sustainability considerations for ethical reasons as ESG-oriented investors, to distinguish them from risk-return investors who care exclusively about the financial performance of a position.
@@
From a theoretical point of view, risk-return ESG does not pose any specific issue, as ESG is treated as any other information (e.g. balance sheet data, macroeconomic indicators, sentiment analysis, etc.) and is integrated in the investment process without affecting the main goal of the investor: improving the risk-adjusted performance. In contrast, the ethical motivations of ESG-oriented investors challenge many of the traditional assumptions adopted in finance theory, as the inclusion of other factors in the investment process (e.g. reduction of CO2 emissions, respect of human right, exclusion of controversial sectors, etc.) are not necessarily motivated by the improvement of financial performance.
@@
To illustrate how the ethical motivation breaks common investing principles, consider two stocks with the same reward–risk profile, but with different ESG scores. They may be equivalent for a risk-return investor but not for an ESG-oriented investor, as one of the two companies may be more environmentally sustainable or may adopt stricter human right policies. An ESG-oriented investor may decide to deliberately worsen their risk-adjusted performance to comply with non-negotiable principles, for instance by excluding certain sectors or companies, thus reducing the diversification of their portfolio. [^3] It is common practice for sustainable investment funds to exclude tobacco, weapons, fossil fuels, and other controversial sectors. The divestment campaign aimed at South Africa’s apartheid regime in the 1980s was one of the key turning points in the history of responsible investing (Schanzenbach and Sitkoff, 2020).
The problem of designing optimal portfolio strategies for ESG-oriented investors that jointly account for financial performance and sustainability considerations has been studied in the portfolio optimization literature. This line of research typically extends the classical reward-risk framework by introducing a third dimension – sustainability – proxied by ESG scores. The scores are incorporated into the portfolio selection problem either as explicit constraints or as additional terms in the objective function. Such an approach has been pursued by, among others, Utz et al. (2015), Gasser et al. (2017) and Cesarone et al. (2022) who studied the efficient frontier of non-dominated portfolios in the reward-risk-ESG space, extending the traditional optimal portfolio literature that started with Markowitz (1952). These works aim to find portfolios with the highest expected return for a given level of risk, where risk is measured using the portfolio variance, a coherent risk measure such as the average value at risk (AVaR) [^4] The AVaR is also known in the literature as Conditional Value at Risk (CVaR) or Expected Shortfall (ES) (Rockafellar et al., 2000), or an asymmetric deviation measure (Giacometti et al., 2021). Pedersen et al. (2021) further contributes to the literature by studying the effect on the market equilibrium: they build a framework with three classes of investors (ESG unaware, that do not consider ESG scores; ESG-aware that use sustainability scores to update their views on risk and expected return; and ESG motivated, that have preferences for high ESG) and they consider an ESG-efficient frontier by considering the Sharpe ratio and the ESG score. The presence of sustainability-oriented investors in the market and they effect on asset prices has been studied also by Heinkel et al. (2001) that investigated the effects of exclusionary ethical investments on corporate behaviour, finding that if the percentage of ethical investors is sufficiently high, the cost of capital for polluting companies may increase. More recently, Pástor et al. (2021) built a two-factor model to study the market equilibrium in the presence of ESG investors, finding that, in equilibrium, sustainable assets have lower expected returns not only due to the fact that they can be used to hedge climate risk, but also because of the preference of ESG investors towards these assets.
One drawback shared by the modeling frameworks discussed so far is that they fail to take into account the stochasticity of ESG scores: using either the expected values of the scores or assuming that ESG scores can be treated as constants. Here, we aim to address this limitation and, more generally, to introduce a framework to jointly consider the financial and ESG dimensions.
In this work we target explicitly ESG-oriented investors, proposing measures that account for the joint distribution of financial returns and ESG scores. In particular, we propose axiomatic classes of measures that extend the concepts of coherent risk measures (Artzner et al., 1999), reward measures (Rachev et al., 2008), and reward-risk ratios (Cheridito and Kromer, 2013). We underline that our framework treats ESG as an explicit argument of investor preferences rather than as a return-predictive signal. Still, the approach can be integrated in asset allocation strategies that jointly incorporate ESG-based preferences and ESG-driven information for financial performance.
@@
Our axiomatic approach is based on the idea that risk comes from two drivers: monetary performance (i.e., the financial returns of a position) and sustainability (represented by the ESG score of the company). [^5] For convenience, in the rest of this paper, we refer to these two dimensions as the monetary and ESG components of risk. These quantities are random and not necessarily independent, and we can represent them as a bivariate random variable. Hence, it is natural to refer to the rich literature on multivariate risk measures (Jouini et al., 2004; Hamel, 2009; Wei and Hu, 2014; Ekeland et al., 2012). Such measures have been developed to study portfolios of non-perfectly fungible assets (e.g., assets valued in multiple currencies) or assets that are difficult to price. We propose using a bivariate risk measure to deal with a single asset that can be evaluated over two dimensions: the monetary returns and the sustainability (represented by ESG scores). We then define ESG-coherent risk measures as an extension of the coherent risk measures introduced by Artzner et al. (1999).
@@
The proposed ESG-coherent measures can be directly embedded in optimal portfolio frameworks (in the objective function or as constraints), with minimal modifications compared to the univariate case. Unlike ESG constraints or screening rules, ESG enters through the risk functional itself, preserving convexity and dual representations. Portfolio choice is therefore formulated as a well-defined optimization problem based on the joint distribution of financial returns and ESG outcomes.
From an applied perspective, the analysis is motivated by market participants’ increasing focus on sustainability and by the emergence of high-frequency ESG measures, such as the Truvalue scores issued by FactSet, as opposed to the largely annual disclosure cycle of conventional ESG data.
Section 2 discusses our interpretation of ESG scores and outlines how we propose to use them. Section 3 introduces ESG-coherent risk measures and provides examples, while Section 4 introduces ESG reward–risk ratios. Section 5 presents an empirical application using the equity returns and real-time ESG data of the constituents of the Dow-Jones Industrial Average (DJIA) financial index. Section 6 discusses policies and practical applications. Section 7 highlights some conclusions.
@@ 2 Measuring ESG and financial performance
The first step in defining ESG risk measures is to clarify how we characterize and measure sustainability. Following an approach common in the literature, we use the ESG scores of a company as a proxy for sustainability (see, e.g., Pedersen et al., 2021; Utz et al., 2015; Cesarone et al., 2022). Here, we clarify how to treat such a variable from both practical and theoretical points of view, establishing the basis for measuring ESG scores in a way that allows us to consistently use them alongside monetary values for risk measurement. We emphasize that, at this stage, we keep the discussion on an abstract level, without reference to specific data providers or methodologies. This allows us not to be limited by the current state-of-the-art market ESG scores, which are still far from being standardized and comparable across data providers (see, e.g., Billio et al., 2021; Berg et al., 2022; Lauria et al., 2026).
@@
Several approaches appear in the literature for modeling the random variable $r_{T}$ used to compute risk. Artzner et al. (1999) define coherent risk measures for a random variable that represents net worth by following the principle that “bygones are bygones,” meaning that future net worth is the only thing that matters. In an alternative approach, the random variable used to compute monetary risk represents the return of a financial position. This latter approach, which is often used in practical applications for the measurement of the risk of equity portfolios, can be interpreted as the payoff of a unitary investment in a risky position, and it introduces some differences in the interpretation of the axioms (see Rachev et al., 2011, Chapter 6). [^6] Alternatively, some authors prefer to define risk measures computed on a variable that represents losses, thus assuming that lower values of the variable are preferred by an investor. Such an assumption does not significantly alter the analysis: it simply changes some signs in the definition of risk (see, e.g., Rockafellar and Uryasev, 2013).
@@
Similarly, we need to establish the quantity measured by the random variable ${ESG}_{T}$ and the sign convention used. In particular, one must consider whether it is a stock variable (measured at a specific point in time) or a flow variable (measured over an interval of time, as some sort of sustainability return). Based on how they are computed, we argue that ESG scores belong to the second category: indeed, they represent the current level of sustainability of the production and commercial practices of a company, which directly affects the impact of that company on the world. [^7] The scores are typically computed as functions of several indicators related to the production methods, the supply chain management, the industry in which the company operates, the transparency of its governance, the presence of specific policies on human rights violations, etc. Unlike static ESG ratings, our interpretation treats sustainability as an accumulated flow of externalities (positive or negative) that are generated over time. In this sense, the ESG score of a company is related to the rate at which it accumulates non-monetary “satisfaction” for the investor. The total non-monetary satisfaction for an investor is proportional to the holding time of the investment.
@@
The variable ${ESG}_{T}$ aggregates sustainability performance over the holding period. We can then interpret the ESG scores assigned by market providers with periodicity $T$ (e.g. a day or a year), as (rescaled) realizations of the random variables ${ESG}_{T}$.
Finally, for each asset we define a bivariate random variable $X_{T}$ used to compute ESG risk and reward measures, as:
$X_{T}=\begin{bmatrix}r_{T}\\ {ESG}_{T}\end{bmatrix}$, (2)
where ${ESG}_{T}$ is the sustainability dimension, and $r_{T}$ is the monetary dimension, computed as the cumulative log-return over a static-horizon of length $T$.
Formally, our framework works under the assumption that the bivariate random vector $X_{T}$ is measurable on the same probability space $(\Omega,\mathcal{F},{P})$. We acknowledge that in practical settings the ESG score process may depend on external information, or that sustainability related information may affect the dynamic of financial returns. We point out that the theoretical axiomatic setup does not impose structural assumptions beyond measurability and joint distribution of $X_{T}$.
The stochasticity of the process $\text{esg}_{t}$ can be traced to two sources: the first is the uncertainty in the measurement of its value, and the second is the uncertainty in the evolution of the sustainability policies within the company, which is driven by the choices made by the management, by market conditions, and by the regulatory framework. We also expect a correlation structure between ${ESG}_{T}$ and financial returns $r_{T}$ due to the presence of common driving factors related, for instance, to sector-wide or country-wide dynamics.
In the rest of the work we use the following scaling convention for the random variable ${ESG}_{T}$: assuming that the ESG scores are bounded, we rescale the values of daily ESG scores such that ${ESG}_{T}\in[-1/c,1/c]$, where $T$ is equal to 1 trading day, and $c=252$. As discussed later, when applied to the Factset Truvalue ESG scores this standardization leads to the “typical magnitude” of ${ESG}_{T}$ for a daily time period to be comparable to the “typical magnitude” of the financial cumulative returns measured over the same time period.
The periodic return $r_{T}$ may be viewed as the payoff of a unit investment in a risky position. Consequently, ${ESG}_{T}$ is taken to refer to the same unit investment, so that the variable is proportional to the initial size of the position. With respect to sign conventions, it is natural to define ESG scores so that higher values are preferable, in line with the standard preference for higher returns. These conventions are adopted for comparability and do not restrict the generality of the axiomatic results.